OpenStudy (anonymous):

WorkPad Note: Remember to show all of the steps that you use to solve the problem for questions 4–6. You can use the comments field to explain your work. Your teacher will review each step of your responses. Given QRS TUV, QS = 3v+2 and TV = 7v – 6, find the length of QS and TV. (2 points)

OpenStudy (anonymous):

The triangles are proportional so you can set 3v+2 equal to 7v-6... 3v+2=7v-6 3v+2+6=7v-6+6 3v+8=7v 3v-3v+8=7v-3v 8=4v v=2 Then plug v into each equation: 3(2)+2=8 and 7(2)-6=8

OpenStudy (anonymous):

did u understand @Rebecca*

OpenStudy (anonymous):

Thank you!!!

OpenStudy (campbell_st):

sorry where does it say they are proportional, or congruent...

OpenStudy (anonymous):

so the length was 8 of both ?